# Multilevel Fast Multipole Algorithm for Electromagnetic Scattering of Metasurfaces using a Static Mode Representation

### v1

@@ IV Fast Multipole Method
We apply the Multilevel Fast Multipole Algorithm (MLFMA) [5, 29, 30, 31] to solve the PMCHWT system (9) using a basis of entire-domain static modes to represent electric and magnetic surface currents. MLFMA accelerates the matrix-vector product by enabling groups of basis functions to interact collectively rather than through pairwise interactions between every possible pair of basis functions.
Initially, basis functions are grouped.
Following this, the impedance matrix $Z$ is decomposed as the sum of two matrices: the near-field and the far-field matrices. The near-field matrix is associated with interactions among basis functions within adjacent groups, while the far-field matrix is associated with interactions between non-adjacent basis functions.
@@
Traditionally, grouping basis functions involves a hierarchical tree structure, such as a Quadtree or Octree [6, 31]. The bounding box, i.e., the smallest box containing the entire object, is first identified. This box is then recursively subdivided until, at the finest level, the size of the box is a fraction of the wavelength, typically $\lambda/4$.
@@
Figure 1: Illustration of the far product in the MLFMA-SME. (a) Particle-particle interaction on level $L_{\max}$. (b) Level $L_{\max}-1$, where $L_{\max}$ is the finest level.
@@
At the finest level, standard Quadtree/Octree approaches might divide a meta-atom into different groups, making the grouping of entire-domain basis functions ambiguous. Leveraging the fact that we deal with meta-atoms whose dimensions are smaller than the wavelength, we propose grouping basis functions by considering each particle as a finest-level group, as shown in Fig. 1(a). At this level, all static modes $\bm{\psi}^{\mu}_{i}$ are physically associated with the center of the particle $\Omega_{\mu}$. For coarser levels, we hierarchically group the basis functions using a standard Quadtree structure (Fig. 1b), which is particularly suitable for structures where the lateral dimensions significantly exceed the vertical ones, as in the case of metasurfaces and metalenses.
@@
Following the fast multipole prescriptions, we decompose the impedance matrix $Z$ as the sum of its block-diagonal part (near part) and the off-diagonal part (far part). For identical particles, the self-particle blocks $Z^{\mu\mu}$ are identical, i.e. $Z^{\mu\mu}=Z_{s}$ for $\mu=1,...,p$. Therefore, $Z_{s}$ can be evaluated and stored once, leading to significant savings in memory requirements and CPU time. To address the singularities in Green’s function when $\mathbf{r}=\mathbf{r}^{\prime}$, the numerical integration of shape function times Green’s function or its gradient are evaluated using the techniques introduced in Refs. [32] and [33].
According to (10), the far-field part of $Z$ is
[…]
with $\mu,\nu=1,...,p$. Since only $k^{-}$ appears when $\mu\neq\nu$, in the following we simplify the notations by omitting the ”--” symbol, thus $k=k^{-}$.
To evaluate the operators $T^{\mu\nu}$ and $K^{\mu\nu}$ when $\Omega_{\mu}$ is far from $\Omega_{\nu}$ (i.e. $\mu\neq\nu$), we rewrite the equations (5) and (6) as
@@
$\frac{e^{-jk\|\mathbf{r}-\mathbf{r}^{\prime}\|}}{4\pi\|\mathbf{r}-\mathbf{r}^{\prime}\|}=\\ \int_{\mathbb{S}^{2}}e^{-jk\hat{\mathbf{k}}\cdot(\mathbf{r}-\mathbf{r}_{\mu})}T_{L}(\hat{\mathbf{k}},\mathbf{r}_{\mu\nu})e^{jk\hat{\mathbf{k}}\cdot(\mathbf{r}^{\prime}-\mathbf{r}_{\nu})}\mathrm{d}\hat{\mathbf{k}},$ (31)
with $\mathbf{r}\in\Omega_{\mu}$, $\mathbf{r}^{\prime}\in\Omega_{\nu}$, $\mathbf{r}_{\mu\nu}=\mathbf{r}_{\mu}-\mathbf{r}_{\nu}$, where $\mathbf{r}_{\mu}$ and $\mathbf{r}_{\nu}$ are, respectively, the centers of the particles $\Omega_{\mu}$ and $\Omega_{\nu}$, $\mathbb{S}^{2}=\{\hat{\mathbf{k}}\in\mathbb{R}^{3}:\|\hat{\mathbf{k}}\|_{2}=1\}$ is the unit sphere, $T_{L}$ is the translation function [6, 34, 35, 31]
@@
where $h_{\ell}^{(2)}$ is the second kind spherical Hankel function of order $\ell$, $P_{\ell}$ is the Legendre Polynomial of order $\ell$.
Using the expansion (31) in Eqs. (28) and (29), we obtain
@@
$K^{\mu\nu}_{mn}=-jk\int_{\mathbb{S}^{2}}\hat{\mathbf{k}}\times\bm{\hat{\psi}}^{\mu}_{m}(\hat{\mathbf{k}})\cdot T_{L}(\hat{\mathbf{k}},\mathbf{r}_{\mu\nu})\bm{\hat{\psi}}^{\nu*}_{n}(\hat{\mathbf{k}})\ \mathrm{d}\hat{\mathbf{k}},$ (34)
where and $\bm{\hat{\psi}}^{\mu}_{m}$ is the $k$-space representation of the basis function
@@
It is well known that truncating the series (32) at a finite order $L$ requires careful handling due to numerical instabilities caused by the nature of spherical Hankel functions [36, 34, 35, 37, 38]. Ideally, one would choose a high order to minimize error, but these functions are highly oscillatory at higher order, meaning that increasing $L$ does not necessarily enhance accuracy [39]. Furthermore, these functions diverge at small arguments, that is, when the electrical distance between two centers is too small, leading to a low-frequency breakdown [34, 39]. However, in this field of application, such instabilities are not observed at optical frequencies, because the particles are not too close together.
Figure 2: (a) First ten longitudinal static modes of a nanodisk. The arrows represent the direction of the surface current density field, the colors represent the surface charge density. (b) First ten transverse static modes of a nanodisk. The arrows represent the direction of the surface current density field, the colors represent the magnitude of the current density field. Modes are presented in ascending lexicographic order and sorted according to their respective static eigenvalues.
@@ V Results and Discussion
Figure 3: (a) Scattering cross section of a finite Vogel spiral composed of 100 gold spheres of radius $R=100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$. $C_{sca}$ is computed using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. (b) Normalized scattered far field at $\lambda$ = 600 nm. The reference Mie solution (dashed line) is also provided for comparison.
The execution of the numerical algorithm can be subdivided into five stages:
  • i)
    Assembly of the self-interaction impedance matrix $Z_{s}$ of an individual meta-atom and the right-hand side vector $\mathbf{V}$ using RWG basis functions. Since the particles are identical, $Z_{s}$ only needs to be evaluated and stored once. The computational complexity and memory requirement for assembling and storing $Z_{s}$, in terms of the number of particles, is $(2n_{e})^{2}\mathcal{O}(1)$.
  • ii)
    Generation of the static modes of an isolated object. Specifically, the longitudinal static modes are obtained by solving the eigenvalue problem (18) using star functions as both expansion and testing functions; the transverse static modes are obtained by solving the eigenvalue problem (21) using loop functions as both expansion and testing functions. In Fig. 2, we present the calculated first ten longitudinal and ten transverse static modes for a nanodisk.
  • iii)
    Evaluation of the MLFMA operators, consisting of the translation function $T_{L}$, the radiation and receive operators $\bm{\hat{\psi}}$ and $\bm{\hat{\psi}}^{\dagger}$. The time complexity and memory requirements for assembling $T_{L}$ are on the order of $\mathcal{O}(p\log p)$. Since the operators $\bm{\hat{\psi}}$ and $\bm{\hat{\psi}}^{\dagger}$ are evaluated for each particle, their time and memory complexities are on the order of $\mathcal{O}(p)$. These complexities can be reduced to $\mathcal{O}(1)$ if the particles have the same spatial orientation (i.e., there is no rotation between particles).
  • iv)
    Compression stage, i.e. transitioning from the representation of the matrices $Z_{s}$, $\bm{\hat{\psi}}$ and $\bm{\hat{\psi}}^{\dagger}$ in terms of the RWG basis to the representation in terms of the static modes basis. This stage can be integrated into steps i) and iii), as there is no need to directly store the RWG representation of these operators.
  • v)
    Solution of the PMCHWT system using an iterative method accelerated by the Multilevel Fast Multipole Algorithm (MLFMA), which reduces the computational complexity of the matrix-vector product from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p\log p)$. This stage is the most time-consuming, as the entire system is solved with a complexity of $\mathcal{O}(Ip\log p)$, where $I$ is the number of iterations required. To accelerate the convergence of the Generalized Minimal Residual Method (GMRES) [40], we employ a block-diagonal preconditioner, where each block is the single-particle self-interaction impedance matrix. We choose a relative tolerance of $10^{-4}$. It effectively reduces $I$ to $\mathcal{O}(p^{\alpha})$, with $\alpha<1$.
@@
In the following simulations, the truncation number $L$ of the translation function (32), is initially set to $10$ at the finest level. This value is subsequently doubled at each successive level [6]. The numerical evaluation of the integrals over the unit sphere is performed using $K=2L^{2}$ point [6]. For interpolation, we employ a $4\times 4$ Lagrange interpolation scheme. The FORTRAN code used to implement this algorithm operates on a single CPU (Intel Xeon-Gold 6140 M 2.3 GHz).
@@ V.1 Validation
We validate the MLFMA-SME method by investigating the scattering from a particle array composed of 100 identical gold nano-spheres, each of radius $R=100$ nm, and arranged according to a golden angle spiral pattern [41]. The $i$-th particle is centered at $(x_{i},y_{i})$, according to
$\begin{cases}x_{i}=r_{i}\cos(i\alpha)\\ y_{i}=r_{i}\sin(i\alpha),\\ \end{cases}$ (36)
where $i=1,...,p$; $r_{i}=c\sqrt{i}$, $c$ is a constant scaling factor, $\alpha=2\pi/\phi^{2}$ is the golden angle, where $\phi=(1+\sqrt{5})/2$ is the golden ratio. We choose the scaling factor as $c=\sqrt{3}R$.
The gold permittivity is obtained by interpolating experimental data [42]. The array is illuminated by a linearly polarized plane, propagating along the axis orthogonal to the array’s plane, linearly polarized along the $x$-axis, with varying wavelength $\lambda$ in the interval $[300,700]$nm. Each particle’s surface is discretized by a triangular mesh having $500$ nodes, $996$ triangles, and $1494$ edges. The scattering cross section, $C_{sca}$, is plotted in Fig. 3 a as function of $\lambda$. The solution, computed by the MLFMA-SME, utilizing $N^{\|}=N^{\perp}=10$ modes per particle, corresponding to a total number of $4000$ unknowns, is compared against the multiparticle Mie theory [19], where the field expansions are truncated at the order $M=15$. The relative error on the scattering cross section is defined as
$\epsilon\{C_{sca}\}=\frac{|C_{sca}-\bar{C}_{sca}|}{\bar{C}_{sca}},$ (37)
where $\bar{C}_{sca}$ is the reference solution obtained by the multiparticle Mie theory. The achieved error is lower than $0.02$ all over the investigated spectral range.
@@
For the same array, in Fig. 3b we show the normalized scattered far field for a fixed wavelength $\lambda=600$ nm. The reference Mie solution, obtained for $M=30$, is included for comparison. An excellent agreement is found.
V.2 Computational costs

### v2

@@ IV Fast Multipole Method
We apply the Multilevel Fast Multipole Algorithm (MLFMA) [17, 53, 54, 55] to solve the PMCHWT system (9) using the static mode basis to represent electric and magnetic surface currents. MLFMA accelerates the matrix-vector product by enabling groups of basis functions to interact collectively rather than through pairwise interactions between every possible pair of basis functions.
@@
Initially, basis functions are grouped. Traditionally, grouping basis functions involves a hierarchical tree structure, such as a Quadtree or Octree [18, 55]. The bounding box, i.e., the smallest box containing the entire object, is first identified. This box is then recursively subdivided until, at the finest level, the size of the box is a fraction of the wavelength, typically $\lambda/4$ [18].
@@
Figure 1: Illustration of the far product in the MLFMA-SMR. (a) Particle-to-particle interaction on the finest level $L_{\max}$. (b) Level $L_{\max}-1$.
@@
At the finest level, standard Quadtree/Octree approaches might divide a meta-atom into different groups, making the grouping of entire-domain basis functions ambiguous. Leveraging the fact that we deal with meta-atoms whose dimensions are smaller than the wavelength, we group basis functions by considering each particle as a finest-level group, as shown in Fig. 1(a). At this level, all static modes $\bm{\psi}^{\mu}_{i}$ are physically associated with the center of the particle $\Omega_{\mu}$. For coarser levels, we hierarchically group the basis functions using a standard Quadtree structure (Fig. 1(b)), which is particularly suitable for structures where the lateral dimensions significantly exceed the vertical ones, as in the case of metasurfaces and metalenses.
@@
Following this, the impedance matrix $Z$ is decomposed as the sum of two matrices: the near-field matrix, and the far-field matrix. The near-field matrix is associated with self-interaction and interactions among basis functions within adjacent groups, while the far-field matrix is associated with interactions between non-adjacent basis functions. When the particles are well separated, namely their distance is greater than $\lambda/2$, only the self-interactions are considered as near. Additionally, for identical particles, the self-particle blocks $Z^{\mu\mu}$ are identical, i.e. $Z^{\mu\mu}=Z_{s}$ for $\mu=1,...,p$. Thus, instead of evaluating and storing $Z^{\mu\nu}$ for all $\Omega_{\nu}$ near to $\Omega_{\mu}$, $Z_{s}$ can be computed and stored once, leading to significant savings in memory requirements and CPU time, compared to classical MLFMA implementation. To address the singularities in Green’s function when $\mathbf{r}=\mathbf{r}^{\prime}$, the numerical integration of shape function times Green’s function or its gradient are evaluated using the techniques introduced in Refs. [56] and [57].
According to (10), only the operators $T_{-}^{\mu\nu}$ and $K_{-}^{\mu\nu}$ appear in $Z^{\mu\nu}$ when $\mu\neq\nu$. Thus, in the following we simplify the notations by omitting the “--” subscript. To efficiently evaluate the $(m,n)$ entry of the operators $T^{\mu\nu}$ and $K^{\mu\nu}$ when $\Omega_{\mu}$ is far from $\Omega_{\nu}$ (i.e. $\mu\neq\nu$), the fast multipole method [17] prescribes to rewrite the matrices in Eq. (11) as
@@
$\frac{e^{-jk\|\mathbf{r}-\mathbf{r}^{\prime}\|}}{4\pi\|\mathbf{r}-\mathbf{r}^{\prime}\|}\approx\int_{\mathbb{S}^{2}}e^{-jk\hat{\mathbf{k}}\cdot(\mathbf{r}-\mathbf{r}_{\mu})}T_{L}(\hat{\mathbf{k}},\mathbf{r}_{\mu\nu})e^{jk\hat{\mathbf{k}}\cdot(\mathbf{r}^{\prime}-\mathbf{r}_{\nu})}\mathrm{d}\hat{\mathbf{k}},$ (26)
with $\mathbf{r}\in\Omega_{\mu}$, $\mathbf{r}^{\prime}\in\Omega_{\nu}$, $\mathbf{r}_{\mu\nu}=\mathbf{r}_{\mu}-\mathbf{r}_{\nu}$, where $\mathbf{r}_{\mu}$ and $\mathbf{r}_{\nu}$ are, respectively, the centers of the particles $\Omega_{\mu}$ and $\Omega_{\nu}$, $\mathbb{S}^{2}=\{\hat{\mathbf{k}}\in\mathbb{R}^{3}:\|\hat{\mathbf{k}}\|_{2}=1\}$ is the unit sphere, $T_{L}$ is the translation function [18, 58, 59, 55] truncated at the order $L$:
@@
where $h_{\ell}^{(2)}$ is the second kind spherical Hankel function of order $\ell$, $P_{\ell}$ is the Legendre Polynomial of order $\ell$. Using the expansion (26) in Eqs. (23) and (24), we obtain
@@
$K^{\mu\nu}_{mn}\approx-jk\int_{\mathbb{S}^{2}}\hat{\mathbf{k}}\times\bm{\hat{\psi}}^{\mu}_{m}(\hat{\mathbf{k}})\cdot T_{L}(\hat{\mathbf{k}},\mathbf{r}_{\mu\nu})\bm{\hat{\psi}}^{\nu*}_{n}(\hat{\mathbf{k}})\ \mathrm{d}\hat{\mathbf{k}},$ (29)
where $\bm{\hat{\psi}}^{\mu}_{m}$ is the $k$-space representation of the basis function
@@
Figure 2: (a) First ten longitudinal static modes of a nanodisk of radius 100 nm and height 50 nm, described by a surface mesh having 588 nodes, 1172 triangles, and 1758 edges. The arrows represent the direction of the surface current density field, the color represents the surface charge density. (b) First ten transverse static modes of a nanodisk. The arrows represent the direction of the surface current density field, the color represents the magnitude of the current density field. Modes are presented in ascending lexicographic order and sorted according to their respective static eigenvalues.
In this context, $\bm{\hat{\psi}}_{m}^{\mu}$ and $\bm{\hat{\psi}}_{m}^{\mu*}$ are also called receiving and radiation patterns, respectively.
The truncation of the translation function (27) at finite order $L$ depends on the electrical size of the groups, $ka$, where $a$ is the group size. An approximation for $L$ is given by [59, 18] $L=kd+1.8d_{0}^{\frac{2}{3}}(kd)^{\frac{2}{3}}$, where $d_{0}$ is the desired number of significant digits, and $d=\sqrt{2}a$. However, truncating the series in Eq. (27) at a finite order $L$ presents challenges due to numerical instabilities caused by the nature of spherical Hankel functions [60, 58, 59, 61]. While increasing $L$ might intuitively seem to reduce error, the oscillatory nature of these functions at higher orders complicates this approach, often leading to diminished accuracy rather than improvement [62]. Furthermore, these functions exhibit divergence at small arguments, corresponding to cases where the electrical distance between two centers is very small, leading to a low-frequency breakdown [58, 62]. Nonetheless, in the context of the investigated metasurfaces, these instabilities are typically absent, as the particles under consideration are sufficiently separated. Furthermore, we can directly assess the maximum error in the Eqs. (28) and (29) as
$\varepsilon_{L}\{A\}=\max_{{\mu,\nu}}\frac{\|A^{\mu\nu}-\bar{A}^{\mu\nu}\|_{2}}{\|\bar{A}^{\mu\nu}\|_{2}}$ (31)
where $\|\cdot\|_{2}$ is the matrix 2-norm, $\bar{A}$ is the matrix $T$ or $K$ evaluated by the Eq. (11), while $A$ is the matrix $T$ or $K$ evaluated by the Eqs. (28)-(29).
@@ V Results and Discussion
The execution of the numerical algorithm can be subdivided into five stages: i) Assembly of the near-field impedance matrix $Z_{N}$, and the right-hand side vector $\mathbf{V}$ using RWG basis functions. ii) Generation of the static modes of an isolated object. Specifically, the longitudinal static modes are obtained by solving the eigenvalue problem (18) using star functions as both expansion and testing functions; the transverse static modes are obtained by solving the eigenvalue problem (21) using loop functions as both expansion and testing functions. As an example, in Fig. 2, we present the calculated first ten longitudinal and ten transverse static modes for a nanodisk of radius 100 nm and height 50 nm. iii) Evaluation of the MLFMA operators, consisting of the translation function $T_{L}$, the receiving and radiation patterns (30). The time complexity and memory requirements for assembling $T_{L}$ are on the order of $\mathcal{O}(p\log p)$. Since the radiation pattern (30) is evaluated for each particle, his time and memory complexities are on the order of $\mathcal{O}(p)$. These complexities can be reduced to $\mathcal{O}(1)$ if the particles have the same spatial orientation (i.e., there is no rotation between particles). iv) Compression stage, i.e.transitioning from the representation of the near matrix and the radiation/receiving patterns in terms of the RWG basis to the representation in terms of the static modes basis. This stage can be integrated into steps i) and iii), as there is no need to directly store the RWG representation of these operators. v) Solution of the PMCHWT system using an iterative method accelerated by the MLFMA, which reduces the computational complexity of the matrix-vector product from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p\log p)$. In particular, we used the Generalized Minimal Residual Method (GMRES) [63]. This stage is the most time-consuming, as the entire system is solved with a complexity of $\mathcal{O}(Ip\log p)$, where $I$ is the number of iterations required. To accelerate the convergence of the GMRES, we employ a block-diagonal preconditioner, where each block is the single-particle self-interaction impedance matrix $Z_{s}$. It effectively reduces $I$ to $\mathcal{O}(p^{\alpha})$, with $\alpha<1$. We choose a relative tolerance of $10^{-4}$.
@@
Figure 3: (a) Scattering cross-section of a finite Vogel spiral composed of 100 gold spheres of radius $R=100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$. $C_{sca}$ is computed using the MLFMA-SMR with $N^{\|}=N^{\perp}=10$. (b) Normalized squared magnitude of the scattered electric field in the far region at $\lambda$ = 600 nm. The reference Mie solution (dashed line) is also provided for comparison. (Inset) Surface mesh used to discretize the meta-atoms in the array, consisting of 500 nodes, 996 triangles, and 1494 edges.
@@
In the following simulations, the truncation number $L$ of the translation function (27), is initially set to $10$ at the finest level. This value is subsequently doubled at each successive level [18]. The numerical evaluation of the integrals over the unit sphere is performed using $K=2L^{2}$ points [18]. For interpolation, we employ a $4\times 4$ Lagrange interpolation scheme. The FORTRAN code used to implement the entire algorithm operates on a single CPU (Intel Xeon-Gold 6140 M 2.3 GHz).
@@ V.1 Validation
We validate the MLFMA-SMR method by investigating the scattering from an array consisting of 100 identical gold nano-spheres, each with a radius $R=100$ nm. The particles are arranged according to a golden angle spiral pattern [64], where the center of the $i$-th particle is located at $(x_{i},y_{i})=r_{i}(\cos(i\alpha),\sin(i\alpha))$, where $i=1,...,p$; $r_{i}=c\sqrt{i}$, $c$ is a constant scaling factor, assumed $c=\sqrt{3}R$, $\alpha=2\pi/\phi^{2}$ is the golden angle, and $\phi=(1+\sqrt{5})/2$ is the golden ratio. This aperiodic pattern is particularly convenient for our study as it is deterministic, and allows us to increase the dimension of the array one particle at a time without appreciably changing the array’s density. The gold permittivity is obtained by interpolating experimental data [65]. The array is illuminated by a linearly polarized plane, propagating along the axis orthogonal to the array’s plane, linearly polarized along the $x$-axis. The wavelength $\lambda$ is varied within the interval $[300,700]$nm. Each particle’s surface is discretized using the triangular mesh shown in the inset of Fig. 3(a).
The scattering cross section $C_{sca}$ is plotted in Fig. 3 as a function of $\lambda$. The solution, computed by the MLFMA-SMR, utilizing $N^{\|}=N^{\perp}=10$ modes per particle (resulting in a total of $4000$ unknowns) is compared against the multiparticle Mie theory [36]. The field $\mathbf{E}_{s}^{(\nu)}$ scattered by the $\nu$-th particle is described in terms of transverse magnetic (TM) $\mathbf{N}_{mn}^{(3)}$, and transverse electric (TE) $\mathbf{M}_{mn}^{(3)}$ vector spherical wave functions [66], which are regular at infinity and have pole at the $\nu$-th particle’s center:
$\mathbf{E}_{s}^{(\nu)}(\mathbf{r})=\sum_{n=1}^{N_{\max}}\sum_{m=-n}^{n}a_{mn}^{(\nu)}\,\mathbf{M}_{mn}^{(3)}(\mathbf{r})+b_{mn}^{(\nu)}\,\mathbf{N}_{mn}^{(3)}(\mathbf{r}),$
where $a_{mn}^{(\nu)}$ and $b_{mn}^{(\nu)}$ are the expansion coefficient. We assume $N_{\max}=15$.
@@
The relative error on the scattering cross section is defined as $\epsilon\{C_{sca}\}={|C_{sca}-\bar{C}_{sca}|}/{\bar{C}_{sca}}$, where $\bar{C}_{sca}$ is the reference solution obtained by the multiparticle Mie theory. The achieved error is lower than $0.02$ all over the investigated spectral range. For the same array, in Fig. 3(b) we show the normalized squared magnitude of the (far) scattered electric field at a given wavelength $\lambda=600$ nm. The reference Mie solution, obtained for $N_{\max}=30$, is included for comparison. An excellent agreement is found.
V.2 Computational time and memory cost